OpenStudy (anonymous):

If the sum of five consecutive odd integers is 115, what is the smallest of the five integers?

OpenStudy (anonymous):

divide 115 by five to find the middle number

OpenStudy (jdoe0001):

if say... we had an odd number like hmmm 7 what would be the "next or consecutive odd" integer?

OpenStudy (anonymous):

let n+1,n+3,n+5,n+7,n+9 be consecutive odd integers whose sum is 115 n+1+n+3+n+5+n+7+n+9=115 5 n+25=115 5 n=115-25=90 n=90/5=18 smallest number is 18+1=19